OpenStudy (anonymous):

Solve the equation using the Zero Product Property. (2x – 4)(2x – 1) = 0

OpenStudy (anonymous):

Recall - the Zero-Product Property means that if: a X b = 0 ; then either a = 0 or b = 0 So you can individually set each factor to 0 and solve. (2x-4)(2x-1)=0 then means: 2x-4= 0 or 2x-1=0 Solve each for x and get your solutions. 2x-4=0 add 4 to both sides 2x=4 divide by 2 on both sides x = 2 2x-1=0 add 1 to both sides 2x=1 divide by 2 on both sides x=1/2 :)